Consider a population consisting of 1, 2, and 3. Suppose samples of size 3 are drawn from this population with and without replacement.


1. Determine the number of sets of all possible samples.



2. List all the possible samples, and compute the mean of each sample.



3. Construct the sampling distribution of the sample means

1
Expert's answer
2022-05-24T18:13:37-0400

We have population values 1,2,3, population size N=3 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{1+2+3}{3}=2"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1+0+1}{3}=\dfrac{2}{3}"


"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{2}{3}}\approx0.8165"

a) Select a random sample of size 3 without replacement. We have a sample distribution of sample mean.


1. The number of possible samples which can be drawn without replacement is "^{N}C_n=^{3}C_3=1."


2.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 1,2,3 & 2 \\
 \hdashline

\end{array}"



3.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 2 & 1 & 2 & 4 \\
 \hdashline

\end{array}"



Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=2=\mu"



The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=4-(2)^2=0= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"\sigma_{\bar{X}}=\sqrt{5.25}\approx2.291288"

b) Select a random sample of size 3 with replacement. We have a sample distribution of sample mean.


1. The number of possible samples which can be drawn with replacement is "N^n=3^3=27."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 1,1,1 & 3/3 \\
 \hdashline
 2 & 1,1,2 & 4/3 \\
 \hdashline
 3 & 1,1,3 & 5/3 \\
 \hdashline
 4 & 1,2,1 & 4/3\\
 \hdashline
 5 & 1,2,2 & 5/3 \\
 \hdashline
 6 & 1,2,3 & 6/3 \\
 \hdashline
 7 & 1,3,1 & 5/3 \\
 \hdashline
 8 & 1,3,2 & 6/3 \\
 \hdashline
 9 & 1,3,3 & 7/3 \\
 \hdashline
 10 & 2,1,1 & 4/3 \\
 \hdashline
 11 & 2,1,2 & 5/3 \\
 \hdashline
 12 & 2,1,3 & 6/3 \\
 \hdashline 
 13 & 2,2,1 & 5/3 \\
 \hdashline 
 14 & 2,2,2 & 6/3 \\
 \hdashline 
 15 & 2,2,3 & 7/3 \\
 \hdashline 
 16 & 2,3,1 & 6/3 \\
 \hdashline 
 17 & 2,3,2 & 7/3 \\
 \hdashline 
 18 & 2,3,3 & 8/3 \\
 \hdashline 
 19 & 3,1,1 & 5/3 \\
 \hdashline 
 20 & 3,1,2 & 6/3 \\
 \hdashline 
 21 & 3,1,3 & 7/3 \\
 \hdashline 
 22 & 3,2,1 & 6/3 \\
 \hdashline 
 23 & 3,2,2 & 7/3 \\
 \hdashline 
 24 & 3,2,3 & 8/3 \\
 \hdashline 
 25 & 3,3,1 & 7/3 \\
 \hdashline 
 26 & 3,3,2 & 8/3 \\
 \hdashline 
 27 & 3,3,3 & 9/3 \\
 \hdashline 
\end{array}"



3.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 3/3 & 1/27 & 3/81 & 9/243 \\
 \hdashline
 4/3 & 3/27 & 12/81 & 48/243 \\
 \hdashline
 5/3 & 6/27 & 30/81 & 150/243 \\
 \hdashline
 6/3 & 7/27 & 42/81 & 252/243 \\
 \hdashline
 7/3 & 6/27 & 42/81 & 294/243 \\
 \hdashline
 8/3 & 3/27 & 24/81 & 192/243 \\
 \hdashline
 9/3 & 1/27 & 9/81 & 81/243 \\
 \hdashline
\end{array}"



Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{162}{81}=2=\mu"



The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{1026}{243}-(2)^2=\dfrac{2}{9}= \dfrac{\sigma^2}{n}"

"\sigma_{\bar{X}}=\sqrt{\dfrac{2}{9}}=\dfrac{\sqrt{2}}{3}\approx0.4714"